OpenStudy (anonymous):

Find the coordinates of the points where C crosses the y-axis, giving your answers as simplified surds if C = (x+2)^2+(y-1)^2=4

OpenStudy (amistre64):

since the yaxis is also the x=0 axis; just make x=0 and solve for y

OpenStudy (anonymous):

do it

OpenStudy (anonymous):

that's the part i got stuck @

OpenStudy (amistre64):

what is 0+2?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

what's 2^2?

OpenStudy (anonymous):

and you tell me what's wrong here

OpenStudy (anonymous):

either the equation is wrong or the exam paper has a mistake in it

OpenStudy (anonymous):

or i'm just stupid

OpenStudy (amistre64):

if the openstudy doesnt boot me out, sure :)

OpenStudy (amistre64):

2^2 is ....

OpenStudy (anonymous):

i just need an answer

OpenStudy (amistre64):

(x+2)^2+(y-1)^2=4 (0+2)^2+(y-1)^2=4 2^2+(y-1)^2=4 im sure you do, but thats not how we play in here

OpenStudy (anonymous):

i can do integration by parts

OpenStudy (anonymous):

the 4s cancel out

OpenStudy (amistre64):

you dont need to integrate by parts for this tho :)

OpenStudy (amistre64):

good, the 4s cancel (y-1)^2 = 0

OpenStudy (anonymous):

so y=1

OpenStudy (anonymous):

but it says answer in surd form?

OpenStudy (amistre64):

y = 1 is what i would get too

OpenStudy (amistre64):

sqrt(1) then

OpenStudy (anonymous):

that would be SR(-1)

OpenStudy (anonymous):

which is i lol

OpenStudy (amistre64):

a surd is just the radical; and i think it means to keep exact answers

OpenStudy (anonymous):

oh wait it's SR1...lmao

OpenStudy (anonymous):

lmaoooo i can't believe i missed that out

OpenStudy (amistre64):

:)

OpenStudy (anonymous):

that's why i hate exams, esp british ones

OpenStudy (amistre64):

the simplified surd would be 1 in my opinion

OpenStudy (anonymous):

why?

OpenStudy (anonymous):

tbh there must be something fishy...how is 1 a surd? >_>

OpenStudy (amistre64):

because a simplified surd can be a rational or integer value sqrt(4) simplifes to 2 sqrt(8) simplifies to 2 sqrt(2)

OpenStudy (amistre64):

its not asking for a surd; its asking for the simplified form of it

OpenStudy (anonymous):

so they're just putting that in to mess with your brains?

OpenStudy (amistre64):

pretty much

OpenStudy (anonymous):

can't you also do y^2+1=0?

OpenStudy (anonymous):

hence y=SR-1

OpenStudy (amistre64):

no (y-1)^2 = 0 y-1 = sqrt(0) = 0 y = 1

OpenStudy (anonymous):

yeah i was stupid for one whole hour...

OpenStudy (anonymous):

btw i got to level57 in one day - yesterday

OpenStudy (anonymous):

is that good?

OpenStudy (amistre64):

dunno, ive never been less than a 99

OpenStudy (anonymous):

so you're a moderator?

OpenStudy (anonymous):

how would one apply for such a title?

OpenStudy (amistre64):

its not an applied for job, its an invitation by the administrators

OpenStudy (anonymous):

is the title bribe-able?

OpenStudy (anonymous):

only kidding - wink wink

OpenStudy (amistre64):

lol, you have to sacrifice chickens; or so ive been led to believe

OpenStudy (anonymous):

:O

OpenStudy (anonymous):

do eggs count?

OpenStudy (anonymous):

chickens are too expensive

OpenStudy (amistre64):

not according to the axiom; never count your chickens b4 they hatch

OpenStudy (anonymous):

haha

OpenStudy (anonymous):

so which really came first? the chicken or the egg?

OpenStudy (anonymous):

that would make a good question on OS biology, if there's such a thing?

OpenStudy (amistre64):

the chicken is the egg, so they have always coexisted; only thing is eggs are uncountable ... imo

OpenStudy (anonymous):

so are you currently studying mathematics @ university?

OpenStudy (anonymous):

are you some kind of wiz professor? :p

OpenStudy (amistre64):

im getting a math degree yes

OpenStudy (anonymous):

can you help me with this?

OpenStudy (anonymous):

∫(x+1)(x+2)(x+3)(x−5)(2+2x3x)(e2x+2x)d

OpenStudy (anonymous):

x

OpenStudy (anonymous):

e^2x btw

OpenStudy (amistre64):

prolly ... but last time there was a division inthere right?

OpenStudy (anonymous):

yesss

OpenStudy (anonymous):

what is the integral of (x+1)(x+2)(x+3)/(x-5)(2+2x^3x)(e^2x+2x) dX from 0 to 1?

OpenStudy (amistre64):

i think your missing some exponential notations ^(..)

OpenStudy (amistre64):

thats better

OpenStudy (amistre64):

..... i come up with -.27691

OpenStudy (anonymous):

how?

OpenStudy (amistre64):

divination of course :)

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

hahahaha^